OpenStudy (anonymous):

What is the remainder when 4^31 is divided by 10?

OpenStudy (anonymous):

its 4

jimthompson5910 (jim_thompson5910):

Hint: 4^1 = 4 (mod 10) 4^2 = 4*4 = 6 (mod 10) 4^3 = 6*4 = 24 = 4 (mod 10) 4^4 = 4*4 = 16 = 6 (mod 10) ... Keep going. I'm sure you can see the pattern

OpenStudy (anonymous):

Thanks a lot it really helped :)